OpenStudy (anonymous):

the sequence an=8,13,18,23, is the same as the sequence a1=8

OpenStudy (anonymous):

an=an1+5

OpenStudy (anonymous):

8,13,18,23 can be defined as a(n) = 5n+3

OpenStudy (anonymous):

im trying to find out if its true or not plz help :"(

OpenStudy (anonymous):

I don't understand what you typed. It is a(n) = n+5?

OpenStudy (anonymous):

its telling me an=an-1+5

OpenStudy (anonymous):

a(2) = a(2)-1+5 a(2) = 13-1+5 a(2) = 12+5 a(2) = 17 Is that what it means? If so, then it is not the same

OpenStudy (anonymous):

thanks